# HumanVLA: Towards Vision-Language Directed Object Rearrangement by Physical Humanoid

### v1

@@ 4 Human-in-the-Room Dataset
Existing datasets [48, 53] for object rearrangement are primarily designed for stretches with disc-shaped mobility and gripper manipulation. Human-like embodiment has different physical attributes, such as stronger motors to handle large furniture like chairs and tables. Besides, we follow [36, 15, 33, 50] to use a humanoid model with spherical hands, which can struggle with manipulating small-sized objects, such as picking up a towel. To address these issues, we introduce a novel Human-in-the-Room (HITR) dataset, designed to facilitate object rearrangement tasks on a humanoid. The HITR dataset includes carefully designed objects of various sizes, ranging from 21cm to 126cm, and provides a variety of rearrangement tasks in various rooms. Each task involves separated, instantiable, and replaceable objects with defined initial and goal states. Additionally, each task is accompanied by a natural language instruction generated by a Large Language Model (LLM). The dataset will be publicly released along with the whole project.
@@ 5.3 Comparisons in Box Loco-Manipulation
InterPhys [15] 60.3 51.4 5.8
HumanVLA-Teacher 80.0 18.9 4.6
Offline GC-BC [29] 11.1 150.9 8.4
HumanVLA 63.7 48.5 5.5
     w/o active rendering 58.8 63.5 5.8
@@ 5.5 Generalizing to Unseen Tasks
We use the test split of HITR to evaluate the generalizability of methods. The unseen data, which includes novel scene compositions, visual appearances, and language instructions, poses a significant challenge to our method. Results are presented in Tab. 4. The state-based HumanVLA-Teacher tends to be more robust in unseen data. Relatively small drops in success rate (5.9%) and precision (4.5cm) demonstrate strong generalizability of RL when using privileged information. Moreover, it consistently outperforms the InterPhys [15] baseline on all metrics. Applying HumanVLA to unseen tasks turns out to be more challenging due to the complexity of vision and language modalities. The success rate of HumanVLA decreases to 63.7%, and the precision drops to 48.5cm. However, HumanVLA still consistently outperforms baselines without active rendering and the offline goal-conditioned behavior-cloning (Offline GC-BC) [29] method.
@@ 5.6 Qualitative Results
We provide qualitative visualizations of how HumanVLA performs object rearrangement tasks in Fig. 4. More results are available in the appendix. Demos are attached in the supplementary material.
@@ D.1 Generation Process
Figure 5 provides an illustration of how we construct the Human-in-the-Room (HITR) dataset, following the procedural generation pipeline [9]. First, we manually design four distinct room layouts: bedroom, livingroom, kitchen, and warehouse. These template layouts are subsequently populated with various object models to create a set of diverse scenes, which serve as the goal states for the rearrangement tasks. To generate the initial state of each task, we randomly relocate an object to a different receptacle or to the ground. The initial position of the humanoid is randomly sampled from the navigable areas. Notably, our dataset does not specify the initial humanoid pose; instead, it is sampled from the training motion dataset. We then concatenate images of the goal and initial states to create a self-contained information carrier for each task. It is sent to a large language model, specifically gpt-4-vision, along with language prompts, to generate the corresponding instruction.
@@ Appendix F Additional Results
F.1 Learning Curves
@@ F.2 Qualitative Ablation
F.3 Additional Qualitative Results

### v2

@@ 4 Human-in-the-Room Dataset
Existing datasets [48, 53] for object rearrangement are primarily designed for stretches with disc-shaped mobility and gripper manipulation. Human-like embodiment has different physical attributes, such as stronger motors to handle large furniture like chairs and tables. Besides, we follow [36, 15, 33, 50] to use a humanoid model with spherical hands, which can struggle with manipulating small-sized objects, such as picking up a towel. To address these issues, we introduce a novel Human-in-the-Room (HITR) dataset, designed to facilitate vision-language directed object rearrangement tasks on a humanoid. The HITR dataset includes carefully designed objects of various sizes, ranging from 21cm to 126cm, and provides a variety of rearrangement tasks in various rooms. Each task involves separated, instantiable, and replaceable objects with defined initial and goal states. Additionally, each task is accompanied by a natural language instruction generated by a Large Language Model (LLM).
@@ 5.3 Comparisons in Box Loco-Manipulation
InterPhys [15] 59.5 52.5 5.9
HumanVLA-Teacher 79.3 19.3 4.6
Offline GC-BC [29] 10.2 152.3 8.5
HumanVLA 60.2 57.0 5.8
     w/o active rendering 56.7 65.5 5.9
@@ 5.5 Generalizing to Unseen Tasks
We use the test split of HITR to evaluate the generalizability of methods. The unseen data, which includes novel scene compositions, visual appearances, and language instructions, poses a significant challenge to our method. Results are presented in Tab. 4. The state-based HumanVLA-Teacher tends to be more robust in unseen data. Relatively small drops in success rate (6.6%) and precision (4.9cm) demonstrate strong generalizability of RL when using privileged information. Moreover, it consistently outperforms the InterPhys [15] baseline on all metrics. Applying HumanVLA to unseen tasks turns out to be more challenging due to the complexity of vision and language modalities. The success rate of HumanVLA decreases to 60.2%, and the precision drops to 57.0cm. However, HumanVLA still consistently outperforms baselines without active rendering and the offline goal-conditioned behavior-cloning (Offline GC-BC) [29] method.
@@ 5.6 Qualitative Results
We provide qualitative visualizations of how HumanVLA performs object rearrangement tasks in Fig. 4. More results are available in the appendix. We demonstrate that HumanVLA is capable of moving a pot, vase, chair, and box based on language instructions.
@@ 6 Conclusion
Acknowledgments
This work was supported by the National Key Research and Development Project of China (No. 2022ZD0160102), National Key Research and Development Project of China (No. 2021ZD0110704), Shanghai Artificial Intelligence Laboratory, and XPLORER PRIZE grants.
@@ Appendix A Limitations
The current version of our system does not contain explicit memorizing, planning, navigation, and multi-agent collaboration modules. We leave more ad-hoc designs to future work.
@@ D.1 Generation Process
Figure 5 provides an illustration of how we construct the Human-in-the-Room (HITR) dataset, following the procedural generation pipeline [9]. First, we manually design four distinct room layouts: bedroom, livingroom, kitchen, and warehouse. These template layouts are subsequently populated with various object models to create a set of diverse scenes, which serve as the goal states for the rearrangement tasks. To generate the initial state of each task, we randomly relocate an object to a different receptacle or to the ground. The initial position of the humanoid is randomly sampled from the navigable areas. As for the initial orientation, the humanoid heads to the object. Our dataset guarantees the object visibility in the first view and also covers the target position visibility in 89% of tasks. Notably, our dataset does not specify the initial humanoid pose; instead, it is sampled from the training motion dataset. We then concatenate images of the goal and initial states to create a self-contained information carrier for each task. It is sent to a large language model, specifically gpt-4-vision, along with language prompts, to generate the corresponding instruction.
@@ Appendix F Additional Results
F.1 Generalization
The primary test set of HITR contains task-level unseen tasks including new compositions of objects in the scene, new placement of objects, and regenerated new text instructions from LLM describing new compositions and new spatial relations. We have proven the generalization ability of HumanVLA in unseen tasks.
However, generalization in any unseen without any similar pattern in the seen data is super challenging, which is also an ultimate goal of embodied AI research. We build extra tiny unseen data to make additional analysis and further disclose our method. We make additional testing data: (1) Unseen texts generated for training tasks, manually reviewed to be distinct from training data. (2) Unseen objects by changing visual appearance in training tasks. (3) Unseen object category (cup) with different geometry. (4) Unseen scene layouts by repositioning static large objects.
Results are reported in Tab. 7. We find that our work suffers less from unseen texts and unseen visual appearance. But generalizing to unseen object categories and execution in the unseen scenes remains a main challenge.
Table 7: Unseen data analysis.
Useen Text 65 50.4 5.4
Unseen object (visual) 50 72.3 6.2
Unseen object (geometry) 20 118.8 7.9
Unseen scene layout 35 88.5 6.8
Figure 12: Additional qualitative results.
F.2 Learning Curves
@@ F.3 Qualitative Ablation
F.4 Additional Qualitative Results